OpenStudy (anonymous):

f(x)=ln(x)+13(x)³-11(x)+cos(x), f'(x)=?

OpenStudy (anonymous):

f'(x) = d(ln(x))/dx + d(13x^3)/dx - d(11x)/dx + d(cosx)/dx

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

ur r8 @sauravshakya

OpenStudy (anonymous):

lets finish it

OpenStudy (anonymous):

Do u know d(lnx)/dx = ??? @jadi

hartnn (hartnn):

\(\huge \frac{d}{dx}\ln x=1/x \\ \huge \frac{d}{dx}x^n=nx^{n-1} \\\huge \frac{d}{dx}\cos x=-\sin x\)

OpenStudy (hba):

@jadi What did you get ?

OpenStudy (anonymous):

looks like he is confused

OpenStudy (hba):

It is easy lol.

OpenStudy (anonymous):

Did u get this? f'(x) = d(ln(x))/dx + d(13x^3)/dx - d(11x)/dx + d(cosx)/dx

OpenStudy (hba):

@jadi Please respond so we can help

OpenStudy (anonymous):

he is sleeping !!1 where is he?

OpenStudy (anonymous):

To find F'(x) we take the derivatives of each i.e f'(x)=(ln(x))'+(13x3)'-(11x)'+(cosx)'-----Look x3 mean x cubic =1/x+13*3x2-11-sinx------look!x2means x squr =26x2-sinx+1/x-11----That's it!!!!

OpenStudy (anonymous):

thanks guys i was just not there but i got it